d Short-Answer Questions 1. U se the quantity equation for this problem. Suppose the money supply is $200, real output is 1,000 units, and the price per unit of output is $1. a. What is the value of velocity? earning. All Rights Reserved. May not be copied, scanned, or duplicated, in whole or in part.
hen inflation gets very high, people do not like to hold money because it is losing value quickly. Therefore, they spend it faster. If when the money supply is doubled, people spend money more quickly, what happens to prices? Do prices more than double, less than double, or exactly double? Why? . W
Answer #1

a) As per Quantity Theory of money (QTM),

MV = PY

Here, P = 1, M= 200 and Y = 1000.

Therefore, V = PY/V = 1*1000/200 = 5

Hence, the velocity of money is 5.

P.S. : As HOMEWORKLIB's policy, if nothing is mentioned than only the first question is to be answered.

Thanks!